A University of Zululand student graduates with a Bachelor of Science degree in Agriculture (Agronomy) by passing 20 modules in total. The marks obtained for each module (in percentage form) are given below: 71 52 81 66 77 96 83 77 65 59 82 97 72 91 50 63 53 74 77 88 1.1 Group the data into a frequency distribution with the lowest class lower limit of 50 percent and a class width of 10 percent. (2 Marks) 1.2 Using the raw data determine the range. (2 Marks) 1.3 Determine the mean and mode from the raw data. (2 Marks) 1.4 Construct an OGIVE curve corresponding to the data. (4 Marks)
1
Expert's answer
2020-04-20T17:57:09-0400

1.1. We build a frequency distribution table with the lowest class lower limit of 50% and a class width of 10%. Then we build a histogram.

1.2. Range=max-min=97-50=47.

1.3. We calculate mean=73.7 and mode=77 of the raw data using AVERAGE and MODE in Excel.

1.4. We build Ogive curve. x-axis is Upper Limits and y-axis is Cumulative Frequency %.